ANTDASAV:

Convert 11.42424242 … to a rational expression in the form of a over b, where b ≠ 0.

Mercury:

this is a little complicated to explain, so please let me know if any step confuses you let's re-write the original expression as 11 + 0.424242... let's let x = the decimal part = 0.424242... we want an expression that puts the repeating part of the decimal on the left side (this is a little easier to demonstrate than explain) if x = 0.424242... then 100x = 42.424242... correct? we just multiplied the original decimal by 100 let's now subtract 100x - x = 99x = 42 (notice how the decimal part "cancels" out with subtraction) if 99x = 42, then x = 42/99. so now that's the original decimal converted to a fraction. so going back to the original expression: 11 + 0.424242 = 11 + 42/99 from there, since they want a rational expression in the form a/b, simply convert this mixed number into an improper fraction